math question

how do you solve (8 + 6a) 4 - 8- = 128?
Answer #1
(8+6a)4-8=128 32+48a-8=128 48a=128-32+8 48a=104 a=104/48= 2.17
answered by: Pooja11
Answer #2

SOLUTION :


(8 + 6a)4 - 8 = 128

Open the bracket :

=> 32 + 24a - 8 = 128

=> (32 - 8) + 24a = 128

=> 24 + 24a = 128

Subtract 24 from both sides :

=> 24a = 128 - 24 

=> 24a = 104 

Divide both sides by 24 :

=> a = 104/24 

=> a = 13/3 = 4.33 (ANSWER).

answered by: Tulsiram Garg